Fermi National Laboratory
 

Accelerator Update for October 9, 2006 - October 11, 2006

 

Monday October 9

The day shift began with Operations monitoring store 5008 (luminosity 71.69E30) and the Recycler stash (134.2E10), with the Antiproton Source (Pbar) stacking (55.4E10), with the Collider Detector at Fermilab (CDF) and D-Zero (D0) using the colliding beam, with Linac, Booster, and the Main Injector (MI) operating normally, and with NuMI and MiniBooNE taking beam.

The Recycler completed stashing antiprotons at 8:34 AM.

The Neutron Therapy Facility (NTF) reported at 10:19 AM that they had completed a patient treatment.

Recycler experts began an hour of reverse proton studies at noon.

The Recycler completed a stash at 1:16 PM.

Pbar experts began an hour of stacktail studies at 1:16 PM.

Operations prepared Pbar for an hour access at 2:41 PM.

NTF reported at 2:50 PM that they had completed their last patient treatment for the day.

Preaccelerator experts accessed the H- Source at 2:54 PM. They came out of access and NuMI and MiniBooNE resumed taking beam at 3:33 PM.

The Recycler completed a stash at 7:47 PM.

Tevatron Electron Lens (TEL) experts began at study period at 9:03 PM.

At 10:31 PM, the Recycler experienced a vacuum burst. Emittances and vacuum were recovering. All vacuum pumps remained on.

The TEL experts ended their study at 11:04 PM.

The Recycler completed a stash at 11:17 PM.

Tuesday October 10

The midnight shift began with Operations monitoring store 5008 (luminosity 33.2E30) and the stash (307.1E10), with Pbar stacking (32.6E10), and with NuMI and MiniBooNE taking beam.

At 1:16 AM, a P1 transfer line power supply (HVF11) tripped off and caused the beam permits for the P1 line and NuMI to trip. Everything reset.

At 1:50 AM, the Recycler lost about 120E10 of its stash. There was a vacuum burst in 30 region; vacuum pumps tripped off causing vacuum valves to close. Operators got everything back on.

At 2:19 AM, Operations reported a power glitch. Store 5008 quenched at Tevatron (TeV) sectors B1 and D1, 2, and 3. The stash and stack survived, but the Linac and Booster went down as well as the RF for the TeV and MI.

At 2:30 AM, Operations reported that they had lost contact with Linac. Operators reported that the Preaccelerator was down along with every Linac RF station. Operations contacted an expert.

At 2:35 AM, Operations reported that D0 was hit hard. Their solenoid was down and they lost all monitoring.

At 2:35 AM, Operations reported that they lost feeder 46B at the master substation. Feeder 46B feeds the cryo system equipment at B0, D0, and F0. Operations contacted the High Voltage Coordinator.

At 2:52 AM, per RF expert instructions, Operations turned on the MI RF system.

NuMI reported at 3:23 AM that everything was okay with their experiment.

At 4:37 AM, the Cryo System Coordinator reported that there might be a bad magnet around the D3 sector. Operators began preparing the TeV for an access.

D0 accessed their collision hall at 5:21 AM.

At 6:36 AM, the High Voltage Coordinator discovered the cause of the problem. A mouse had managed to crawl through a cable entry and short the switchgear.


If you are missing your pet mouse, please come to the MCR to claim him and please bring proof of ownership.

At 7:01 AM, a TeV expert reported that he found a possible ground fault.

Operations reported at 7:29 AM that Linac controls had returned.

At 8:51 AM, Operations reported that experts had found a bus to bus to ground fault. Experts continued their investigations.

At 10:09 AM, Operations reported that a NuMI LCW water skid had faulted. Experts were contacted.

TeV experts isolated D2 and D3 and successfully hipotted the rest of the TeV at 12:36 PM. The TeV D3 magnet faulted. It will take about a week to replace.

At 1:37 PM, Preaccelerator experts reported that the H- Source was shorted. They are working to clear the short. If they can’t, it will take about twenty-four hours before the I- source system would be ready.

Linac experts reported at 4:04 PM that they had completed their LRF2 repair.

Operations prepared Pbar at 7 PM for an extraction study. The study ended at 8:17 PM.

A Pbar expert began a study at 8:50 PM.

Wednesday October 11

The midnight shift began with Operations monitoring the stash (177.5E10) and the Pbar stack (19.1E10), with Pbar and Recycler experts conducting studies, with the TeV off for a magnet replacement, and with all other beam off due to problems with the H- Source (the I- Source will be brought online during the day shift).

The latest Pbar study ended at 4:04 AM. Operations then prepared the area for an access.

The Recycler studies ended at 7 AM.

Experts began work on the MI-10 kickers at 7:21 AM.

The Plans for Wednesday and the rest of the Week

  1. Replace the Tevatron D3 magnet
  2. Check the Tevatron separator vacuum at sectors D0 and D17 – the TeV separators may require a bake
  3. Fix the Pbar magnet LCW leak (A6Q14)
  4. Fix the NuMI magnet LCW leak (I:HV100) – this may require a magnet change
  5. Allow the Pelletron experts to conduct necessary maintenance

Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now.

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.

last modified 10/11/2006    email Fermilab